If the probabilities are 0.87, 0.36, and 0.29 that, while under warranty, a new car will require repairs in the engine, drive train, or both, what is the probability that a car will require one or the other or both kinds of repairs under the warranty?


1
Expert's answer
2021-03-31T16:54:08-0400

"P(A) = 0.87 \\

P(B) = 0.36 \\

P(A \cap B) = 0.29"

We have to find:

"P(A \cup B) = P(A) + P(B) - P(A \cup B) \\

= 0.87 + 0.36 -0.29 \\

= 0.94"